OpenStudy (anonymous):

solve for x. show all work and all the steps. 5x + 7 = 42

OpenStudy (anonymous):

1-subtract 7 in each side : 5x+7-7=42-7 5x=35 2-divide by 5 each side : 5x/5=35/5 x=7

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

how about this, 3x - 7 + 4x = 42